Via Gelato is a popular neighborhood gelato shop. The company has provided the following cost formulas and actual results for the month of June:

Fixed Element
per Month
Variable Element
per Liter
Actual Total
for June
Revenue $ 16.00 $ 90,540
Raw materials $ 5.05 $ 31,630
Wages $ 6,000 $ 1.80 $ 16,900
Utilities $ 2,030 $ 0.60 $ 6,000
Rent $ 3,000 $ 3,000
Insurance $ 1,750 $ 1,750
Miscellaneous $ 690 $ 0.75 $ 5,200

While gelato is sold by the cone or cup, the shop measures its activity in terms of the total number of liters of gelato sold. For example, wages should be $6,000 plus $1.80 per liter of gelato sold and the actual wages for June were $16,900. Via Gelato expected to sell 6,000 liters in June, but actually sold 6,200 liters.

Required:

Calculate Via Gelato revenue and spending variances for June. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1
Flexible budget Actual Revenue and spending variances
Revenue 99200 90540 8660 U
Expenses:
Raw materials 31310 31630 320 U
Wages 17160 16900 260 F
Utilities 5750 6000 250 U
Rent 3000 3000 0 U
Insurance 1750 1750 0 U
Miscellaneous 5340 5200 140 F
Total expense 64310 64480 170 U
Net Operating income 34890 26060 8830 U
Workings:
Flexible budget
Revenue =6200*16
Expenses:
Raw materials =6200*5.05
Wages =6000+(6200*1.8)
Utilities =2030+(6200*0.6)
Rent 3000
Insurance 1750
Miscellaneous =690+(6200*0.75)